OpenStudy (anonymous):

igor invested $2,176 in a saving accounts which earns 3,14% interest compounded quarterly. what will the account be worth in seven years??? A) $2,759.85 B) $21,502.71 c) $2,309.22 d) $5,564.29 please help me

OpenStudy (anonymous):

\[2176\times \left(1+\frac{.0314}{4}\right)^{28}\] and a calculator

OpenStudy (anonymous):

the 28 because four times seven is 28